# On the CP Nature of the ‘95 GeV’ Anomalies

### v2

@@ 2.2 Allowed Parameter Space
In order to explain all the anomalies, the scalar $S$ must be coupled to third-generation fermions as well as massive gauge bosons. We fix the CPV parameter $\alpha$ and scan the free parameters $\rho_{f}^{S}$ to satisfy the excesses defined in Eq. (1), Eq. (4) and Eq. (5). We have shown the parameter space in Fig. 1, which can simultaneously explain all the excesses. For a CP-even scalar, the cross-section is relatively smaller, as shown in Eq. (8), and consequently, a large value of $\rho_{t}$ is required, as depicted in the top left panel of Fig. 1. For a mixed CP state, the required values of $\rho_{t}$ decreases as the cross-section increases due to a pseudoscalar component, the parameter is shown in the top-right panel of Fig. 1. In the bottom panel, we have considered a pure pseudoscalar model. Note that such a pure pseudoscalar scenario can not explain the LEP $b\bar{b}$ excess. Thus, the $g_{SVV}$ is redundant, and the parameter space does not depend on $g_{SVV}$. In all cases, the colour represents the $\rho_{b}$ value. From the figure, it is evident that it is possible to satisfy all three excesses simultaneously.
If we exclude the LEP excess, we can take $g_{SVV}=0$. Such a choice alters the decay spectrum of $S$ as $S\to Zff$ is not possible. In addition, it is possible to assume $\rho_{b}^{S}=0$ to explain the LHC excesses. Hence, to explain the LHC anomalies, as given in Eq. (1)-(4), a lower value of $\rho_{t/\tau}^{S}$ compared to the parameter space shown in Fig 1 will be enough, which corroborates Iguro:2022dok . Note that in Fig. 1, the parameter space does not depend on the sign of the angle $\alpha$, and thus, the results are identical for both ±$\pm\alpha$. As discussed in the next section, the CP observables will distinguish $+\alpha$ from $-\alpha$.
@@ 3 Uncovering the CP Nature of the Resonance
The CP transformation property of a scalar particle can be inferred from its decay to tau pairs since the correlation of the tau lepton polarisation planes is governed by the CP properties of the parent particle. From the reconstruction of the decay planes of the tau leptons, it is possible to infer the CP properties of the scalar particle.
@@
The differential decay width of the scalar boson can be written as Kramer:1993jn ; Berge:2014sra ; Berge:2014wta
$d\Gamma_{S\to\tau^{+}\tau^{-}}\propto 1-\dfrac{\pi^{2}}{16}b(E_{+})b(E_{-})\cos(\phi_{CP}-2\alpha),$ (10)
@@
where $b(E_{\pm})$ are the spectral function describing the spin analyzing power of a given decay mode Tsai:1971vv ; Berge:2014wta . For $\tau\to\pi\nu$ decay, $b(E_{\pm})=1$, and for other decay channels, it depends on the energy of the charged hadron at their respective tau-lepton rest frame. There are several methods  Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua to estimate $\phi_{CP}$ and in this work we have used Impact Parameter (IP) method Berge:2014sra ; Berge:2015nua and the $\rho$ decay plane method  Bower:2002zx ; Berge:2015nua ; Han:2016bvf ; Dutta:2021del for hadronic $\tau$ decay. For completeness, we have described the methods briefly.
3.1 Impact Parameter Method
In the Impact Parameter (IP) method Berge:2014sra ; Berge:2015nua , the $\tau$-lepton decay plane is constructed using the 3-momentum vector ($\vec{q}^{\pm}$) and the three-dimensional (3D) impact parameter vector ($\vec{n}_{\pm}$) of the charged particles. The method can be used for any one-prong decay of the tau lepton, and in this work we have implemented the method for $S\to\tau^{+}\tau^{-}\to(\pi^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decay. The IP vector in the laboratory frame is defined as the distance of the closest approach of the charged particle’s ($\pi^{\pm}$) track to the reconstructed primary vertex (PV) of an event. It is possible to reconstruct the 3D IP vector using the distance between the point of closest approach and the primary vertex in the $x-y$-plane ($d_{0}$) and the $z$-coordinate of the point ($z_{0}$ansen:2020ecn . We have used $d_{0}$ and $z_{0}$ and charged track momentum (highest $p_{T}$ charged track inside a $\tau$-tagged jet) provided by Delphes to find out the IP vector using the procedure describe in Hansen:2020ecn . The IP vector then converted to a four vector $n_{\pm}^{\mu}=(0,\vec{n}_{\pm})^{T}$ and boosted to zero-momentum frame (ZMF), the $\pi^{+}\pi^{-}$ rest frame. The boosted IP four-vectors are denoted as $n_{\pm}^{*\mu}$ and we define the angle,
$\phi^{*}=\arccos(\vec{n}^{*}_{+_{\perp}}.~{}\vec{n}^{*}_{-_{\perp}}),$ (11)
@@ 3.1 Impact Parameter Method
where $n^{*\mu}_{\pm_{\perp}}$ is the perpendicular component of $n^{*\mu}_{\pm}$ along the charged track $h^{\pm}$. To obtain a signed angle, we define the triple product, $\mathcal{O}(=q^{*-}\cdot(\vec{n}^{*}_{+_{\perp}}.~{}\vec{n}^{*}_{-_{\perp}}))$ where $q^{*-}$ is the 3-momentum vector of $\pi^{-}$ in ZMF. The CP observable is given by,
$\phi_{CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&\mathcal{O}\geq 0\\ &\tilde{\phi}^{*}\quad&\mathcal{O}<0.\end{aligned}\right.$ (12)
3.2 Rho Decay Plane Method
@@
When the $\tau$ lepton decays via the rho meson, $\tau\to\rho^{\pm}\nu_{\tau}\to\pi^{\pm}\pi^{0}\nu_{\tau}$, we use the information from the charged and neutral pion to define a CP-odd observable. In the $\pi^{+}\pi^{-}$ rest frame, the acoplanarity angle $\phi^{*}$ is defined between the plane spanned by the $(\pi^{+},\pi^{0})$ and $(\pi^{-},\pi^{0})$ vectors. The acoplanarity angle is given by Bower:2002zx ; Berge:2015nua ; Hansen:2020ecn ,
$\phi^{*}=\arccos(\vec{q}^{~{}*0}_{+_{\perp}}.~{}\vec{q}^{~{}*0}_{-_{\perp}}),$ (13)
@@
where $\vec{q}^{~{}*0}_{+(-)_{\perp}}$ is the 3-momentum of the $\pi^{0}$ associated with $\pi^{+(-)}$ in the $\pi^{+}\pi^{-}$ rest-frame transverse to the direction of the associated charged pion. The signed angle $\tilde{\phi}^{*}$ is obtained using the CP-odd triple correlation $\mathcal{O_{CP}}(=q^{*-}\cdot(q^{*0}_{+_{\perp}}\times~{}q^{*0}_{-_{\perp}})$) and we define
$\tilde{\phi}^{*}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O_{CP}}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O_{CP}}<0.\end{aligned}\right.$ (14)
@@
In the $S\to\tau^{+}\tau^{-}\to(\pi^{+}\pi^{0}\nu_{\tau})(\pi^{-}\pi^{0}\bar{\nu}_{\tau})$ decay, the spin analyzing power is proportional to $y_{\rho}=\dfrac{(E_{\pi^{+}}-E_{\pi^{0}})}{(E_{\pi^{+}}+E_{\pi^{0}})}\times\dfrac{(E_{\pi^{-}}-E_{\pi^{0}})}{(E_{\pi^{-}}+E_{\pi^{0}})}$ at the respective $\tau^{\pm}$ rest frame Bower:2002zx . Since $y_{\rho}$ can be both positive and negative and if we integrate over the charged and neutral pion moments, $\phi_{CP}$ does not display $CP$ sensitivity. We categorize the events depending on $y_{\rho}$ and define the final $CP$ sensitive variable,
@@
$\phi_{CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&y_{\rho}<0\\ &\tilde{\phi}^{*}\quad&\textrm{ otherwise}.\end{aligned}\right.$ (15)
@@
To estimate $y_{\rho}$, we have calculated the energies of the pions in the lab frame, as it is not possible to reconstruct the $\tau$ rest frames. The choice of lab frame energy does not significantly reduce the asymmetry in $\phi_{CP}$ as long as the $\rho$ mesons are energetic ($p_{T}(\rho^{\pm})>20$ GeV) Berge:2015nua which is in general true for LHC.
3.3 Combined IP and Rho Decay Plane Methods
@@
In the case of $\tau^{+}\tau^{-}\to(\rho^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decay, the angle is defined by Berge:2015nua ,
$\phi^{*}=\arccos(\vec{q}^{~{}*0}_{+_{\perp}}.~{}\vec{n}^{~{}*}_{-_{\perp}}),$ (16)
where, $\vec{q}^{~{}*}_{\pm}$ is the 3-momentum of $\pi^{\pm}$ and $\vec{n}^{~{}*}_{-}$ is the IP vector of $\pi^{-}$ and tall the vectors are in the $\pi^{+}\pi^{-}$ rest frame. As before, we can define a signed angle depending on the value of the triple product $\mathcal{O}_{3}(=\vec{q}^{~{}*}_{-}\cdot(\vec{q}^{~{}*0}_{+_{\perp}}\times~{}\vec{n}^{~{}*}_{-_{\perp}})$).
On the other hand, if $\tau^{+}$ decay to $\pi^{+}\nu_{\tau}$, we can define $\phi^{*}$ and the triple product appropriately. Note that, in both cases of mixed decay, we have considered the event classification based on $y_{\pm}(=(E_{\pi^{\pm}}-E_{\pi^{0}}))$.
@@ 3.4 Event Simulation and Selection
We implemented our simplified model for event simulation in FeynRules-2.3Alloul:2013bka . The signal and the background events are generated using MadGraph5-aMC@NLO-3.5.3Alwall:2011uj ; Alwall:2014hca . To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-3.5.3. For showering and hadronization we used PYTHIA-8.3 Sjostrand:2014zea ; Bierlich:2022pfr , and detector effects were simulated using Delphes-3.5.0deFavereau:2013fsa . We use anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV. For $\tau$-tagging we have used medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency. The signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$. We veto events with any $b$-tagged jets.
The background for such signal at the LHC is dominated by $Z\to\tau^{+}\tau^{-}$, $W^{\pm}+jets$ and QCD multi-jet events. We have generated sample background data to understand the nature of BG distribution for $CP$ sensitive observable. As we will discuss later, we have used actual data provided by the CMS collaboration through HEPDATA Maguire:2017ypu and rescaled it accordingly for different luminosity.
We obtain the charged and neutral pions from the $\tau$-jet constituents. From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion. Neutral pions decay promptly into a pair of photons and therefore deposit their energies in the electromagnetic calorimeter, ECAL. From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion. For the combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau$ ($\tau\to\pi\nu_{\tau}$).

### v3

@@ 2.2 Allowed Parameter Space
In order to explain all the anomalies, the scalar $S$ must be coupled to third-generation fermions as well as massive gauge bosons. We fix the CPV parameter $\alpha$ and scan the free parameters $\rho_{f}^{S}$ to satisfy the excesses defined in Eq. (1), Eq. (4) and Eq. (5). We have shown the parameter space that can simultaneously explain all the excesses in Fig. 1. For a CP-even scalar, the cross section is relatively small, as shown in Eq. (8), and consequently, a large value of $\rho_{t}$ is required, as depicted in the top-left panel of Fig. 1. For a CP-mixed state, the required values of $\rho_{t}$ decrease as the cross section increases due to a pseudoscalar component, the parameter is shown in the top-right panel of Fig. 1. In the bottom panel, we have considered a pure pseudoscalar model. Note that such a pure pseudoscalar scenario can not explain the LEP $b\bar{b}$ excess. Thus, the $g_{SVV}$ is redundant, and the viable parameter space does not depend on it. In all cases, the color represents the $\rho_{b}$ value. From the figure, it is evident that it is possible to satisfy all three excesses simultaneously.
If we exclude the LEP excess, we can take $g_{SVV}=0$. Such a choice alters the decay spectrum of $S$ as $S\to Zff$ is not possible. In addition, it is possible to assume $\rho_{b}^{S}=0$ to explain the LHC excesses. Hence, to explain the LHC anomalies, as given in Eqs. (1)–(4), a lower value of $\rho_{t/\tau}^{S}$ compared to the parameter space shown in Fig 1 will be enough, which corroborates Iguro:2022dok . Note that, in Fig. 1, the parameter space does not depend on the sign of the angle $\alpha$, and thus, the results are identical for both $\pm\alpha$.
@@ 3 Uncovering the CP Nature of the Resonance
The CP transformation property of a scalar particle can be inferred from its decay to tau pairs since the correlation of the tau lepton polarization planes is governed by the CP properties of the parent particle. From the reconstruction of the decay planes of the tau leptons, it is then possible to infer the CP properties of the scalar particle.
3.1 Hadronic $\tau^{\pm}$ decays
@@
The differential decay width of a spin-0 boson into two fermions can be written as Kramer:1993jn ; Berge:2014sra ; Berge:2014wta
$d\Gamma_{S\to\tau^{+}\tau^{-}}\propto 1-\dfrac{\pi^{2}}{16}b(E_{+})b(E_{-})\cos(\phi_{\rm CP}-2\alpha),$ (10)
@@
where $b(E_{\pm})$ are the spectral function describing the spin analyzing power of a given decay mode Tsai:1971vv ; Berge:2014wta . For $\tau^{\pm}\to\pi^{\pm}\nu$ decays, $b(E_{\pm})=1$ while, for other decay channels, it depends on the energy of the charged hadron at the respective tau-lepton rest frame. There are several methods  Bower:2002zx ; Desch:2003rw ; Berge:2013jra ; Berge:2008wi ; Berge:2008dr ; Berge:2011ij ; Berge:2014sra ; Berge:2014wta ; Berge:2015nua to estimate $\phi_{\rm CP}$ and in this work we have used the Impact Parameter (IP) one Berge:2014sra ; Berge:2015nua and the $\rho$ decay plane one Bower:2002zx ; Berge:2015nua ; Han:2016bvf ; Dutta:2021del , for hadronic $\tau$ decays. For completeness, we will briefly describe the two methods next.
3.1.1 Impact Parameter Method
[…]
$\phi^{*}=\arccos(\hat{n}^{*}_{+_{\perp}}.~\hat{n}^{*}_{-_{\perp}}),$ (11)
@@ 3.1.1 Impact Parameter Method
where $\hat{n}^{*}_{\pm_{\perp}}$ is the normalized perpendicular component of $\vec{n}^{*}_{\pm}$ along the charged track $\pi^{\pm}$. To obtain a signed angle, we define the triple product, $\mathcal{O}(=\hat{q}^{*}_{-}\cdot(\hat{n}^{*}_{+_{\perp}}.~\hat{n}^{*}_{-_{\perp}}))$ where $\hat{q}^{*}_{-}$ is the normalized 3-momentum vector of $\pi^{-}$ in the ZMF. The CP observable is given by
$\phi_{\rm CP}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O}<0.\end{aligned}\right$. (12)
3.1.2 Rho Decay Plane Method
@@
When the $\tau$ lepton decays via the rho meson, $\tau^{\pm}\to\rho^{\pm}\nu_{\tau}\to\pi^{\pm}\pi^{0}\nu_{\tau}$, we use the information from the charged and neutral pion to define a CP-odd observable. In the $\pi^{+}\pi^{-}$ rest frame, the acoplanarity angle $\phi^{*}$ is defined between the plane spanned by the $(\pi^{+},\pi^{0})$ and $(\pi^{-},\pi^{0})$ vectors. The acoplanarity angle is given by Bower:2002zx ; Berge:2015nua ; Hansen:2020ecn ,
$\phi^{*}=\arccos(\hat{q}^{~*0}_{+_{\perp}}.~\hat{q}^{~*0}_{-_{\perp}}),$ (13)
@@
where $\hat{q}^{~*0}_{\pm_{\perp}}$ is the normalized 3-momentum of the $\pi^{0}$ associated with $\pi^{\pm}$ in the $\pi^{+}\pi^{-}$ rest-frame transverse to the direction of the associated charged pion. The signed angle $\tilde{\phi}^{*}$ is obtained using the CP-odd triple correlation $\mathcal{O_{\rm CP}}(=\hat{q}^{*}_{-}\cdot(\hat{q}^{*0}_{+_{\perp}}\times~\hat{q}^{*0}_{-_{\perp}})$) and we define
$\tilde{\phi}^{*}=\left\{\begin{aligned} &\phi^{*}\quad&\mathcal{O_{\rm CP}}\geq 0\\ &2\pi-\phi^{*}\quad&\mathcal{O_{\rm CP}}<0.\end{aligned}\right$. (14)
@@
In the $S\to\tau^{+}\tau^{-}\to(\pi^{+}\pi^{0}\nu_{\tau})(\pi^{-}\pi^{0}\bar{\nu}_{\tau})$ decay, the spin analyzing power is proportional to $y_{\rho}=\dfrac{(E_{\pi^{+}}-E_{\pi^{0}})}{(E_{\pi^{+}}+E_{\pi^{0}})}\times\dfrac{(E_{\pi^{-}}-E_{\pi^{0}})}{(E_{\pi^{-}}+E_{\pi^{0}})}$ at the respective $\tau^{\pm}$ rest frame Bower:2002zx . Since $y_{\rho}$ can be both positive and negative and if we integrate over the charged and neutral pion moments, $\phi_{\rm CP}$ does not display CP sensitivity. We categorize the events depending on $y_{\rho}$ and define the final CP-sensitive variable,
@@
$\phi_{\rm CP}=\left\{\begin{aligned} &\tilde{\phi}^{*}+\pi\quad&y_{\rho}<0\\ &\tilde{\phi}^{*}\quad&\textrm{ otherwise}.\end{aligned}\right$. (15)
@@
To estimate $y_{\rho}$, we have calculated the energies of the pions in the lab frame, as it is not possible to reconstruct the $\tau$ rest frames. The choice of lab frame energy does not significantly reduce the asymmetry in $\phi_{\rm CP}$ as long as the $\rho$ mesons are energetic ($p_{T}(\rho^{\pm})>20$ GeV) Berge:2015nua which is in general true for LHC.
3.1.3 Combined IP and Rho Decay Plane Methods
@@
In the case of $\tau^{+}\tau^{-}\to(\rho^{+}\nu_{\tau})(\pi^{-}\bar{\nu}_{\tau})$ decays, the angle is defined by Berge:2015nua ,
$\phi^{*}=\arccos(\hat{q}^{~*0}_{+_{\perp}}.~\hat{n}^{~*}_{-_{\perp}}),$ (16)
where, $\hat{q}^{~*0}_{+}$ is the normalized 3-momentum of $\pi^{0}$ (originating from $\rho^{+}$) and $\hat{n}^{~*}_{-}$ is the IP vector of $\pi^{-}$ and all the vectors are in the $\pi^{+}\pi^{-}$ rest frame. As before, we can define a signed angle depending on the value of the triple product $\mathcal{O}_{3}(=\hat{q}^{~*}_{-}\cdot(\hat{q}^{~*0}_{+_{\perp}}\times~\hat{n}^{~*}_{-_{\perp}})$). Instead, if $\tau^{+}$ decays to $\pi^{+}\nu_{\tau}$, we can define $\phi^{*}$ and the triple product appropriately. Note that, in both cases of mixed decay, we have considered the event classification based on $y^{\rho}_{\pm}\Big{(}=\dfrac{E_{\pi^{\pm}}-E_{\pi^{0}}}{E_{\pi^{\pm}}+E_{\pi^{0}}}\Big{)}$.
3.2 Semi-leptonic Decays
For the semi-leptonic channels, $S\to\tau^{+}\tau^{-}\to(\ell^{+}\nu_{\ell}\bar{\nu}\tau)(\tau_{h}^{-}\nu\tau)$ (and charge conjugate), where the leptons denote electrons and muons, the hadronic $\tau$ decay is reconstructed using either the IP method or the $\rho$ decay plane method, as described earlier. The leptonic side is treated as a one-prong decay, and we are using the impact parameter of the charged lepton for our analysis.
[…]
@@ 3.3 Event Simulation and Selection
We implemented our simplified model for event simulation in FeynRules-2.3 Alloul:2013bka . The signal and the background events are generated using MadGraph5-aMC@NLO-3.5.3 Alwall:2011uj ; Alwall:2014hca . To handle the $\tau$ polarization, we have used the TauDecay Hagiwara:2012vz package in MadGraph5-aMC@NLO-3.5.3. For showering and hadronization we used PYTHIA-8.3 Sjostrand:2014zea ; Bierlich:2022pfr while detector effects were simulated using Delphes-3.5.0 deFavereau:2013fsa . We use the anti-kt algorithm Cacciari:2008gp for jet clustering with radius parameter $R=0.5$ and $p_{T}(j)>20$ GeV. For $\tau$-tagging, we have used a medium tag point with $70\%$ tagging and $5\times 10^{-3}$ mis-tagging efficiency.
We follow Ref. CMS:2022goy for event selection. For the hadronic channel, the signal events are classified as two $\tau$-tagged jets with $p_{T}>40$ GeV and $|\eta|<2.1$. For the semi-leptonic events, electrons (muons) should have $|\eta|<2.1$, and need to satisfy isolation criterion of $I_{\text{rel}}<0.15$. The electrons (muons) with $p_{T}>$ 25 GeV (20) GeV is required for the $e-\tau_{h}$ ($\mu-\tau_{h}$) mode. The hadronically decaying tau ($\tau_{h}$) candidate is required to have $|\eta|<2.3$. A $p_{T}$ threshold of 35 GeV (32 GeV) is applied when the event is selected by an $e-\tau_{h}$ ($\mu-\tau_{h}$) pair modes. For both the hadronic and the semi-leptonoc modes, We veto events with any $b$-tagged jets, and we have considered events having $p_{T}^{\tau^{+}\tau^{-}}>200$ GeV as shown in the right panel of Fig. 9 in Ref. CMS:2022goy , where $p_{T}^{\tau^{+}\tau^{-}}$ is obtained by the vectorial $p_{T}$ sum of the visible decay products of tau leptons. Such $p_{T}^{\tau^{+}\tau^{-}}$ criteria help to substantially minimize the background events.
We obtain the charged and neutral pions from the $\tau$-jet constituents. From all the tracks within each $\tau$-jet, the track with the highest $p_{T}$ is considered to be the charged pion. Neutral pions decay promptly into a pair of photons and therefore deposit their energies in the electromagnetic calorimeter. From all the energy deposits, called towers, we estimate the one with the highest $E_{T}$ and assign the tower to the neutral pion. For the IP-$\rho$ combination method, the charged pion track with higher $p_{T}$ is associated with the two-body decay of the $\tau^{\pm}$ ($\tau^{\pm}\to\pi^{\pm}\nu_{\tau}$). In the case of the semi-leptonic mode, the track associated with the lepton is considered while, for the hadronic part, we have used either the IP or the $\rho$-decay method.
The background for such signal at the LHC is dominated by $Z\to\tau^{+}\tau^{-}$, $W$ + jets, $t\bar{t}$ and QCD multi-jet events. We have generated sample background data to understand the nature of the corresponding distribution for our CP-sensitive observable. As we will discuss later, we have used actual data provided by the CMS collaboration through HEPDATA Maguire:2017ypu ; hepdata.128147 and rescaled the background accordingly for a different luminosity.